OpenStudy (anonymous):

(x^2 x^3)^2 .....how do I simplify that?

OpenStudy (anonymous):

(x^2x^3)^2=(x^(2+3))^2 =(x^5)^2=x^10 Remember when two terms are multiplied their powers are added.When a term already raised to a power is further raised to a power the two powers are multiplied

OpenStudy (anonymous):

(X^2 X^3)^2 =X^4 X^6 =X^10